# ThinK: Thinner Key Cache by Query-Driven Pruning

### v1

@@ 4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and KV compression methods integrated with our channel pruning method of Key cache (ThinK) over two different base LLMs across various KV-sizes on LongBench. The following observations can be drawn: (1) Our method can further prune the channels of the key cache after compressing the KV cache with H2O and SnapKV. For the base model LLaMA3-8B, our method reduces memory usage and slightly improves performance for both H2O and SnapKV. For the base model Mistral-7B, our method reduces memory with only a slight drop in performance in some cases. (2) Comparing the results of SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm shown in Table 1, our proposed query-based channel pruning shows better performance when the pruning ratio is $40\%$ compared to no pruning. However, $l_{1}$ and $l_{2}$ norm methods cannot outperform the KV compression methods without pruning. (3) When the pruning ratio is small, performance tends to be better than when the pruning ratio is large. (4) When the KV-size is increased from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV cache size of $2048$ and a pruning ratio of $40\%$, our method can even outperform the LLaMA3-8B with a full KV cache. The above observations indicate that our method is agnostic to existing KV cache compression methods and can further improve their performance and memory reduction. Additionally, query-based pruning is more effective than $l_{1}$ and $l_{2}$ norm for channel pruning in LLMs.
@@ 4.3 Results on Needle-in-a-Haystack
Table 4 presents the results of the Needle-in-a-Haystack test based on the SnapKV (Li et al.,, 2024) approach with KV cache size of input prompt from $128$ to $2048$. With a modest pruning ratio $\lambda=40$%, the accuracy of ThinK is consistently better or comparable than the original SnapKV across both LLaMA3 and Mistral with different KV cache sizes. Such comparisons demonstrate that the proposed query-driven channel pruning can effectively retain the informative channels and discard the noisy ones. With a pruning ratio of $\lambda\geq 50$%, we observe an accuracy drop with a small KV-cache size, particularly for $128$ and $512$ across both LLaMA3 and Mistral. However, ThinK is still able to obtain comparable performance with SnapKV when we have a larger KV cache size (i.e., $1024$ and $2048$). Intuitively, a larger pruning ratio with a smaller KV cache size could tend to lose more important information compared to a larger KV cache size. In addition, the performance on larger KV cache size also suggests ThinK is robust for long context retrieval tasks.
Figure 6 (a)-(d) visualize the Needle-in-a-Haystack test accuracy against different token lengths and depth percent. The KV cache size is configured to $128$ and $1024$, with pruning ratio $\lambda$ equal to $40$% and $50$%, respectively. First, ThinK can maintain the original retrieval capability of SnapKV though the overall numerical accuracy is different (i.e., $77.8$ vs $78.6$ and $90.4$ vs $90.6$). Because ThinKachieves the same accuracy at most of the entries with different token limits and depths as Snap KV. In addition, ThinK also successfully retrieves some “needles” where SnapKV fails, leading to higher overall accuracy. The visualization results further demonstrate the robustness of ThinK in a fine-grained perspective and ThinK can enhance the original quantization approach.
@@
LLaMA3-8B SnapKV 00 79.6 90.2 91.2 91.7
SnapKV+ThinK 40 79.6 90.3 91.2 91.7
SnapKV+ThinK 50 77.4 89.6 91.0 91.7
Mistral-7B SnapKV 00 77.8 89.5 90.4 90.8
SnapKV+ThinK 40 78.6 90.1 90.6 90.9
SnapKV+ThinK 50 78.1 90.1 90.8 91.1
SnapKV+ThinK 60 75.9 89.2 90.6 91.1
@@ 4.4 Ablation Studies
Table 5: Performance comparison of pruning K cache using different recent-size.
@@
SnapKV + ThinK($\lambda=40\%$)
@@
Table 6: ablations with the same memory consumption.
@@
Preserve the most recent KV embeddings (Zhang et al., 2024c, ; Li et al.,, 2024) is important for maintaining the performance of LLMs after KV cache compression. Note that a tradeoff exists: increasing the recent-size allows more infomation to be propagated, while increasing the cache size to be stored. To study its impacts, we evaluate the performance produced by three recent-size, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the basleline model. The results are summarized in Table 5. One can observe that a recent-size of 32 yields superior performance than 00 in terms of averaged score on LongBench which demonstrates the importance of keeping the most recent KVs. On the other hand, the performance of $32$ and $128$ is negligible, suggesting that maintaining the most recent 32 KVs suffices to preserve optimal performance.
Figure 5: Performance comparisons under the same KV cache memory consumption.
Comparison Under the Same Memory Consumption.
@@
To ensure a fair comparison, we adjust the KV cache size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK. For example, the KV cache size of H2O with ThinK is $128$. Due to channel pruning of the key cache, the memory consumption of H2O with ThinK at a $128$ KV-size is lower than H2O at a 128 KV-size. Therefore, the KV-size of H2O is adjusted from $128$ to $109$. Table 6 and Figure 5 present the results of these comparisons on LongBench. As the results show, H2O or SnapKV with ThinK deliver superior performance compared to H2O or SnapKV without ThinK while maintaining the same memory usage. This demonstrates that KV cache compression methods, combined with our query-based channel pruning, can compress the KV cache more effectively and utilize memory more efficiently.
Pruning Channels of Both Key and Value Cache.
@@
In our previous study, we evaluate methods with pruning channels of the key cache. In this part, we explore the impact of pruning channels of the value cache (Appendix A.1). Specifically, for KV cache compression methods, we prune channels of the key cache at one pruning ratio and channels of the value cache at another pruning ratio. Table 7 presents the results with LLaMA3-8B and Mistral-7B on LongBench. When testing on the base model LLaMA3-8B, H2O or SnapKV with key and value channel pruning can perform on par with H2O or SnapKV without channel pruning. In some cases, H2O or SnapKV with key and value channel pruning may even perform better. Pruning the channels of the value cache over the base model Mistral-7B experiences a slight performance drop. Despite this, note that KV cache with additional value channel pruning can further reduce memory usage.
@@
Table 7: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench. H2O $+$ ThinK($\lambda_{1}$+$\lambda_{2}$) indicates that H2O has $\lambda_{1}$of the Key cache channels pruned and $\lambda_{2}$of the Value cache channels pruned.
@@
+ThinK(30+30) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
+ThinK(40+40) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
Mistral 7B KV-size 128
@@
+ThinK(30+30) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
+ThinK(30+10) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94
(a) KV-size 128, SnapKV, Acc.77.8
(b) KV-size 128, SnapKV+ThinK($40$) Acc.78.6
(c) KV-size 1024, SnapKV, Acc.90.4
(d) KV-size 1024, SnapKV+ThinK($50$), Acc.90.8
Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2.
5 Related Work
In scenarios involving long contexts, the most significant computational burden from the attention mechanism is the key-value (KV) cache. Reducing the KV cache is a high priority for optimizing deployment efficiency. System-level optimizations, such as FlashAttention (Dao,, 2023) and PagedAttention (Kwon et al.,, 2023), have been developed to address this issue. Additionally, algorithm-level optimizations are being explored to further enhance efficiency.
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b, ) maintains a few initial tokens and some recent tokens based on the observation of attention sink, which may result in the loss of important information carried by the dropped tokens. H2O (Zhang et al., 2024c, ) retains only a small portion of the tokens by greedily dropping tokens based on their contributions to the cumulative attention. SnapKV (Li et al.,, 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al.,, 2023) adaptively evicts tokens from attention heads that emphasize local contexts. This approach focuses on discarding non-special tokens centered on special tokens, while the standard KV cache is used only for attention heads that broadly attend to all tokens. PyramidKV (Zhang et al., 2024b, ) and PyramidInfer (Yang et al.,, 2024) considers adjusting the KV cache size across different layers by allocating more cache in the lower layers and less in the higher ones.
KV Cache Quantization. SmoothQuant (Xiao et al., 2023a, ) can quantize the KV cache to 8-bit with minimal performance degradation. Q-Hitter (Zhang et al., 2024c, ) uses accumulated attention scores and ‘Quantization Friendliness’ metrics to identify tokens that are essential for maintaining the generalization capabilities of LLMs and are suitable for KV cache quantization. Some studies have found that the key cache and value cache should be quantized differently (Liu et al., 2024b, ; Hooper et al.,, 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
Structured Pruning of LLMs. Structured pruning  (Ma et al.,, 2023; Ding et al.,, 2023) of LLMs traditionally focuses on removing unimportant layers, heads, and hidden dimensions, which often results in significant performance degradation. In contrast, our methodology preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels based on data dependant criterion, our approach greatly reduce the key cache-size with negligible performance loss.
6 End Note and Future Direction
Motivated by the observation that certain channels have significantly larger magnitudes compared to others, and the singular value analysis indicates that the key cache is inherently low-rank (Singhania et al.,, 2024), we propose ThinK to perform pruning over the key cache channel. The proposed pruning strategy is query-dependant and optimized based on the attention scores, ensuring that essential information is retained for each input query. In addition, ThinK can be seamlessly integrated with other popular token-level KV cache quantization techniques (Li et al.,, 2024; Zhang et al., 2024c, ), further enhancing inference efficiency. Extensive experiments on LongBench (Bai et al.,, 2023) and Needle-in-a-Haystack tests with two foundation models demonstrate the effectiveness and robustness of our query-dependent channel pruning method. Our approach achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%. Our analysis indicates that ThinK can maintain superior performance over baselines with a smaller KV cache size under equivalent memory consumption conditions.
Future work will focus on strategies to increase the pruning ratio without compromising model performance, and exploring pruning techniques for value caches. We also plan to evaluate channel importance from a more sophisticated and compositional perspective, incorporating both token-level and channel-level information to calculate importance.
Appendix A Additional Experimental Details
A.1 Value Cache Pruning
$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (11)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache. Then, only top T channels are retained.

### v2

@@ 4.2 Results on LongBench
Tables 2 and  3 present the results of KV compression methods and their integration with our proposed channel pruning technique for the key cache (ThinK) across three different base LLMs, evaluated at various KV-sizes on the LongBench benchmark. The following observations can be drawn: (1) Our method successfully prunes the channels of the key cache after the KV cache has been compressed using H2O and SnapKV. For the LLaMA-3-8B-Instruct base model, our approach reduces memory usage while slightly improving performance for both H2O and SnapKV. For the Mistral-7B-Instruct-v0.2 base model, our method similarly reduces memory usage, with only a minor performance drop in some cases. For the larger LLaMA-3-70B base model, our method achieves comparable or superior performance after pruning $40\%$ of the key cache channels, compared to the SnapKV baselines. (2) When Comparing SnapKV or H2O integrated with ThinK in Table 2 to SnapKV or H2O integrated with $l_{1}$ or $l_{2}$ norm in Table 1, our query-driven channel pruning approach demonstrates superior performance when the pruning ratio of $\lambda=0.4$. (3) Lower pruning ratios generally result in better performance compared to higher pruning ratios. (4) As the KV-size increases from $128$ to $2048$, the performance of our channel pruning method improves. Notably, with a KV-size of $2048$ and a pruning ratio of $0.4$, our method even surpasses the performance of LLaMA-3-8B-Instruct with a full KV cache. These findings suggest that our method is agnostic to the underlying KV cache compression techniques and can further enhance both performance and memory efficiency. Moreover, query-driven channel pruning proves more effective than $l_{1}$ and $l_{2}$ norm-based methods for channel pruning in LLMs.
We further validate the efficacy of our method by applying it to the KV cache quantization technique KIVI (Liu et al., 2024b), as shown in Table 4. First, we prune $40\%$ of the key cache channels, followed by quantization of the remaining channels into 2-bit (implementation details provided in Appendix C.1). Compared to the standard KIVI method, our approach reduces KV cache memory by $20\%$, with minimal performance degradation.
@@ 4.3 Results on Needle-in-a-Haystack
LLaMA3-8B-Instruct SnapKV 0.0 79.6 90.2 91.2 91.7
SnapKV+ThinK 0.4 79.6 90.3 91.2 91.7
SnapKV+ThinK 0.5 77.4 89.6 91.0 91.7
Mistral-7B-Instruct-v0.2 SnapKV 0.0 77.8 89.5 90.4 90.8
SnapKV+ThinK 0.4 78.6 90.1 90.6 90.9
SnapKV+ThinK 0.5 78.1 90.1 90.8 91.1
SnapKV+ThinK 0.6 75.9 89.2 90.6 91.1
Table 5 presents the results of the Needle-in-a-Haystack test, using the SnapKV (Li et al., 2024) approach with varying KV-sizes, ranging from $128$ to $2048$. With a modest pruning ratio of $\lambda=0.4$, ThinK consistently outperforms or matches the accuracy of the original SnapKV across both LLaMA-3 and Mistral models, regardless of KV-size. These comparisons demonstrate that the proposed query-driven channel pruning method effectively retains informative channels while discarding noisy ones. However, when the pruning ratio increases to $\lambda\geq 0.5$, we observe a drop in accuracy with smaller KV-sizes, particularly for $128$ and $512$, across both LLaMA-3 and Mistral models. Despite this, ThinK achieves comparable performance with SnapKV when the KV-size is larger(i.e., $1024$ and $2048$). Intuitively, a larger pruning ratio with a smaller KV-size may lead to the loss of more critical information compared to scenarios with a larger KV-size. In addition, the performance on larger KV-sizes suggests that ThinK is robust for long-context tasks.
Figure 6 (a)-(d) (in Appendix B) visualize the Needle-in-a-Haystack test accuracy across varying token lengths and depth percentages. The KV-sizes are set to $128$ and $1024$, with pruning ratios of $\lambda=0.4$ and $\lambda=0.5$, respectively. ThinK preserves the retrieval capabilities of SnapKV, although there are minor numerical differences in overall accuracy (e.g., $77.8$ vs. $78.6$ and $90.4$ vs. $90.6$). ThinK matches SnapKV in accuracy for the majority of token limits and depths, demonstrating consistency in performance. Furthermore, ThinK successfully retrieves certain ”needles” that SnapKV fails to capture, resulting in improved overall accuracy. These visualizations highlight the robustness of ThinK from a fine-grained perspective, illustrating its capacity to enhance the original approach.
@@ 4.4 Ablation Studies
Table 6: Performance comparison of key cache pruning with varying recent-sizes.
@@
SnapKV + ThinK ($\lambda=0.4$)
@@
Table 7: Performance comparison of key cache pruning with the same memory consumption.
@@
Impact of Different Recent Sizes. Preserving the most recent KV embeddings (Zhang et al., 2024c; Li et al., 2024) is important for maintaining the performance of LLMs after KV cache compression. However, a tradeoff exists: increasing the recent-size allows more information to be retained, but also increases the cache size. To assess its impact, we evaluate the performance of three recent-size configurations, namely 00, $32$ and $128$, on LongBench, using Mistral-7B-Instruct-v0.2 as the baseline model. The results are summarized in Table 6. As observed, a recent-size of 32 yields superior performance compared to 00, as indicated by the averaged score on LongBench, demonstrating the importance of retaining the most recent KVs. On the other hand, the performance difference between recent-sizes of $32$ and $128$ is negligible, suggesting that retaining the most recent 32 KVs is sufficient to maintain optimal performance.
Figure 3: (a) presents the performance comparison with token eviction methods under identical memory usage for Mistral-7B-Instruct-v0.2, while (b) illustrates the memory usage comparison with the KV cache quantization method KIVI across different batch sizes for LLaMA-2-7B-chat. ThinK ($0.4$) indicates we prune the key cache channels with a pruning ratio of $\lambda=0.4$.
@@
Performance Comparison Under the Same Memory Usage. To ensure a fair comparison, we adjust the KV-size of H2O or SnapKV to match the memory usage of H2O with ThinK or SnapKV with ThinK on Mistral-7B-Instruct-v0.2. For example, the KV-size of H2O with ThinK is set to $128$. Due to channel pruning applied to the key cache, the memory consumption of H2O with ThinK at a KV-size of $128$ is lower than that of H2O at the same KV-size. Consequently, the KV-size of H2O is adjusted from $128$ to $109$ to equalize memory usage. Table 7 and Figure 3(a) present the results of these comparisons on the LongBench benchmark. The results demonstrate that H2O or SnapKV combined with ThinK consistently outperforms their counterparts without ThinK while maintaining the same memory footprint. This highlights the effectiveness of integrating query-driven channel pruning with KV cache compression methods, enabling more efficient memory utilization and improved compression of the KV cache.
Memory Usage Comparison. To evaluate the efficiency of ThinK, we follow the methodology used in KIVI (Liu et al., 2024b). We generate synthetic workloads with an input prompt length of 160 and an output length of 338. The peak memory usage is reported for the vanilla FP16 baseline, KIVI, and KIVI combined with ThinK ($0.4$) for LLaMA-2-7B-chat. As in Figure 3(b), the memory savings from our method become increasingly evident as the batch size grows, in both the KIVI $/2$ 2 and KIVI $/4$ 4 configurations. Compared to the baseline model, our approach achieves over a 5$\times$ increase in batch size while maintaining the same memory footprint when integrated with KIVI.
@@
Pruning Channels of Both Key and Value Cache. In our previous study, we evaluate methods that focused on pruning channels of the key cache. In this part, we explore the impact of pruning channels in the value cache as well (Appendix D). Specifically, for KV cache compression methods, we apply different pruning ratios to the channels of the key and value caches. Table 8 presents the results with LLaMA-3-8B-Instruct and Mistral-7B-Instruct-v0.2 on the LongBench benchmark. When evaluating the base model LLaMA-3-8B-Instruct, H2O or SnapKV with both key and value channel pruning perform comparably to H2O or SnapKV without pruning. In certain instances, the models with key and value channel pruning even outperform their non-pruned counterparts. For the base model Mistral-7B-Instruct-v0.2, pruning the value cache channels leads to a slight performance drop. Nevertheless, it is important to note that applying value channel pruning to the KV cache further reduces memory usage.
5 Related Work
In scenarios involving long contexts, the key-value (KV) cache poses the most significant computational and memory burden within the attention mechanism of large language models. Reducing the KV cache is therefore a key priority for optimizing deployment efficiency. To address this challenge, system-level optimizations, such as FlashAttention (Dao, 2023) and PagedAttention (Kwon et al., 2023), have been developed to tackle this challenge. In parallel, algorithm-level optimizations are also being explored to further enhance efficiency.
KV Cache Eviction. StreamingLLM (Xiao et al., 2023b) retains a few initial tokens along with some recent tokens based on the observation of attention sinks, which can leading to the loss of critical information carried by the dropped tokens. H2O (Zhang et al., 2024c) selectively retains a small subset of tokens by greedily dropping those with lower contributions to cumulative attention. SnapKV (Li et al., 2024) selects clustered important KV positions for each attention head from an ‘observation’ window located at the end of the prompts. FastGen (Ge et al., 2023) adaptively evicts tokens from attention heads that focus on local contexts, discarding non-special tokens that surround key tokens, while standard KV cache is applied to attention heads that attend more broadly. PyramidKV (Zhang et al., 2024b) and PyramidInfer (Yang et al., 2024) take a hierarchical approach, adjusting KV cache sizes across different layers by allocating more cache to lower layers and less to higher ones.
KV Cache Quantization. SmoothQuant (Xiao et al., 2023a) enables the quantization of the KV cache to $8$-bit with minimal performance degradation. Q-Hitter (Zhang et al., 2024c) leverages accumulated attention scores and ”Quantization Friendliness” metrics to identify tokens that are crucial for preserving the generalization capabilities of LLMs, making them suitable for KV cache quantization. Furthermore, recent studies suggest that the key and value caches should be quantized differently (Liu et al., 2024b; Hooper et al., 2024): the key cache should be quantized per-channel, while the value cache should be quantized per-token.
Structured Pruning of LLMs. Traditional structured pruning  (Ma et al., 2023; Ding et al., 2023) of LLMs typically focuses on removing unimportant layers, heads, or hidden dimensions, often leading to significant performance degradation. In contrast, our approach preserves the original architecture of the LLM and specifically targets the channel dimension within each head’s key cache. By dynamically identifying unimportant channels using data dependant criterion, our method greatly reduce the key cache size with minimal performance loss.
6 Conclusion
Inspired by the observation that certain channels exhibit significantly larger magnitudes than others, and supported by singular value analysis indicating that the key cache is inherently low-rank, we propose ThinK  as a pruning method targeting the key cache channels. Our pruning strategy is query-dependent and optimized based on attention scores, ensuring that essential information is preserved for each input query. In addition, ThinK can be seamlessly integrated with existing token-level KV cache pruning techniques (Li et al., 2024; Zhang et al., 2024c) and KV cache quantization methods (Liu et al., 2024b), further enhancing inference efficiency. Extensive experiments on LongBench and Needle-in-a-Haystack benchmarks demonstrate the effectiveness our query-dependent channel pruning approach. Our method achieves comparable or superior performance to baseline methods while reducing the key cache size by $40$%. For instance, ThinK integrated with KIVI can achieve a 2.8×2.8$\times$2.8 × reduction in peak memory usage while maintaining nearly the same quality, enabling up to a 5×5$\times$5 × increase in batch size when using a single GPU.
Appendix A Observations
Figure 5 and Figure 4 illustrates the observations which motivates our approach ThinK to prune unimportant key cache channels. We conducted singular value decomposition (SVD) (Demmel, 1997) on the attention weights of the specified layers to investigate their principal components. Note that
$\mathbf{U},\mathbf{\Sigma},\mathbf{V}=\text{SVD}\bigg{(}\text{softmax}\Big{(}\frac{\textbf{Q}\textbf{K}^{T}}{\sqrt{D}}\Big{)}\bigg{)},\quad\text{Energy}_{i}=\frac{\sigma_{i}^{2}}{\sum_{i}\sigma_{i}^{2}}.$
(a) Key Cache ($14^{th}$ layer)
(b) Value Cache ($14^{th}$ layer)
(c) Key Cache ($20^{th}$ layer)
(d) Value Cache ($20^{th}$ layer)
Figure 4: Magnitude of key and value cache for LLaMA-2-7B. The first head of layer $14$ and layer $20$ of LLaMA-2-7B is selected to visualize the magnitude of the key and value caches. We observe that the magnitudes of the key cache channels vary differently, whereas the channels of the value cache do not exhibit such variation.
(a) Energy of singular values
(b) Cumulative energy of singular values
Figure 5: The energy and cumulative energy of the singular values.
Appendix B Needle-in-a-Haystack test performance comparison
Figure 6 visualizes the test performance comparison on Needle-in-a-Haystack on Mistral-7B-Instruct-v0.2.
(a) KV-size $128$, SnapKV, Acc. $77.8$
(b) KV-size $128$, SnapKV + ThinK ($0.4$) Acc. $78.6$
(c) KV-size $1024$, SnapKV, Acc. $90.4$
(d) KV-size $1024$, SnapKV + ThinK ($0.5$), Acc. $90.8$
Figure 6: Needle-in-a-Haystack test performance comparison with Mistral-7B-Instruct-v0.2. ThinK ($\lambda$) indicates we prune the key cache channels with a pruning ratio of $\lambda$
Appendix C Implementations
C.1 Implementation with quantization
Figure 7 illustrates the implementation of our method when integrated with the KV cache quantization method KIVI (Liu et al., 2024b). During the prefill phase, we first prune the unimportant channels of $X_{K}$ before applying quantization. In the decoding phase, each newly arrived key cache $t_{K}$ is added to $X_{K_{r}}$. Once $X_{K_{r}}$ reaches $G$ tokens, the residual length hyperparameter in KIVI, we prune and quantize the data, then concatenate it with the previously quantized $Q(P(X_{K_{g}}))$.
Figure 7: Implementations of ThinK when incorporated with KIVI.
Appendix D Value Cache Pruning
$I_{i}=\textbf{Top}_{T}(\text{Score}_{v,i},T)$ (4)
where $\textbf{Q}_{i},\textbf{K}_{i},\textbf{V}_{i}\in\mathbb{R}^{S\times D}$. We define a criterion $\text{Score}_{v,i}$ to indicate the importance of each channel in the head $i$ of value cache. Then, only top T channels are retained. Table 8 reported the results of pruning both key and value channels.
@@ Appendix D Value Cache Pruning
Table 8: Performance comparison of pruning both K and V cache with different pruning ratios on LongBench. H2O $+$ ThinKV ($\lambda_{1}$+$\lambda_{2}$) indicates that the key cache channels of H2O are pruned with a pruning ratio of $\lambda_{1}$ and the value cache channels are pruned of a pruning ratio of $\lambda_{2}$.
@@
+ThinKV (0.3+0.3) 25.13 29.97 40.35 44.12 34.64 19.94 23.62 23.03 25.30 72.50 90.78 39.46 5.35 69.00 61.50 57.91 41.41
+ThinKV (0.4+0.4) 25.13 28.85 40.70 44.21 36.36 21.07 22.31 22.89 24.80 72.50 90.88 38.77 6.41 69.00 61.49 58.87 41.52
Mistral-7B-Instruct-v0.2 KV-size 128
@@
+ThinKV (0.3+0.3) 27.01 30.72 48.81 41.15 26.93 18.93 25.81 23.59 26.42 70.00 86.82 41.91 3.05 82.65 57.01 54.25 41.57
+ThinKV (0.3+0.1) 26.22 32.69 48.96 40.83 26.70 19.02 27.87 24.23 26.64 70.00 86.65 42.63 2.22 85.13 57.00 54.28 41.94